OpenStudy (anonymous):

1. 12s+13=25 ? 2.r/-7 -5=-6 ? 3. 7k-8k=1 ? 4.10-d=19 ? 5. 9/3-4+9=7 ? 6. 5b+6-6b +2=19?

OpenStudy (anonymous):

1) 12s +13 = 25 12s = 12 s = 1 2) r/-7 -5=-6 r/-7 = -1 r = 7 3) 7k-8k=1 -k = 1 k = -1 4) 10-d=19 10-19 = d -9 = d 6) 5b+6-6b +2=19 -b +8 = 19 -b = 11 b = 11 Check 5th one

OpenStudy (anonymous):

a/3-4+9=7